OpenStudy (anonymous):

How can you solve the equation 8x = 2x + 2 graphically?

OpenStudy (anonymous):

@Lopez______RvB

OpenStudy (anonymous):

well x = 1/3

OpenStudy (anonymous):

are there choices

OpenStudy (anonymous):

nope :C

OpenStudy (anonymous):

that was all i knew actually :P

OpenStudy (anonymous):

is the answer supposed to be a number or a written sentence

OpenStudy (anonymous):

its supposed to be typed and the teacher reads it and thats it :l no multiple choice sadly

OpenStudy (anonymous):

o wait um i did a mistake!

OpenStudy (anonymous):

its 8x=2^x+2!!!

OpenStudy (anonymous):

8x = 2x + 2 In order to bring this equation to graphable form, we need to move all the terms to the left side. In our example, - 2 terms , , will be moved to the left side. Notice that a term changes sign when it 'moves' from one side of the equation to the other.

OpenStudy (anonymous):

wait but i messed up cause i didnt realize it look at it again :P

OpenStudy (anonymous):

8x=2^x+2!!!

OpenStudy (anonymous):

that was totally my fault!

OpenStudy (anonymous):

We need to get rid of expression parentheses. If there is a negative sign in front of it, each term within the expression changes sign. Otherwise, the expression remains unchanged. In our example, there are no negative expressions

OpenStudy (anonymous):

but um lopez i gave you the wrong equation

OpenStudy (anonymous):

In order to solve this non-linear equation, we need to move all the terms to the left side. In our example, - 2 terms 2x^x, 2, will be moved to the left side. Notice that a term changes sign when it 'moves' from one side of the equation to the other

OpenStudy (anonymous):

what i just wrote is the correct one

OpenStudy (anonymous):

just write what i wrote it should work

OpenStudy (anonymous):

could i tell you what i think tho so you can explain it for me a little?

OpenStudy (anonymous):

i thought of just grapghing both of them and where they cross is the answer :P

OpenStudy (anonymous):

i am not certain but i would go with what you think is correct

OpenStudy (anonymous):

o ok :D

OpenStudy (anonymous):

good luck